# Federated Distributed Key Generation

### v2

@@ V Security Analysis
In this section, we provide a formal security analysis of the Federated Distributed Key Generation (FDKG) protocol ($\Pi_{FDKG}$). We adopt the standard simulation-based security framework commonly used for Distributed Key Generation (DKG) and Multi-Party Computation (MPC) protocols [15, 30], addressing the need for rigorous definitions and proofs aligned with cryptographic literature. This framework involves comparing the execution of the real protocol in the presence of an adversary with an ideal execution involving a simulator and an ideal functionality.
V-A Security Model
  • Parties: We consider a set of $n$ potential parties $\mathbb{P}=\{P_{1},\dots,P_{n}\}$.
  • Communication Model: Parties communicate via a secure broadcast channel (emulating a bulletin board where messages are public and authenticated). Communication occurs in synchronous rounds.
  • Adversary: We consider a static Probabilistic Polynomial-Time (PPT) adversary $\mathcal{A}$ who corrupts a subset $\mathcal{C}\subset\mathbb{P}$ of parties at the beginning of the protocol execution. The adversary controls the corrupted parties entirely, receiving their inputs and determining their messages. Let $|\mathcal{C}|$ be the number of corrupted parties.
  • Participation: Participation in the FDKG generation phase is optional. The set of actual participants $\mathbb{D}\subseteq\mathbb{P}$ emerges during the protocol execution.
  • Setup Requirements: The protocol requires a pre-existing Public Key Infrastructure (PKI) where each party $P_{i}$ has published an authenticated public encryption key $\textrm{pk}_{i}$ (with corresponding secret key $\textrm{sk}_{i}$) for the private channel mechanism (Section III-F). Additionally, the use of NIZK (specifically Groth16 in our implementation) presupposes the existence of a Common Reference String (CRS), crs, generated via a trusted setup procedure or an equivalent setup mechanism.
  • Cryptographic Assumptions: The security proofs rely on the standard cryptographic assumptions associated with the underlying primitives:
    • PKE Security: The public-key encryption scheme $(\texttt{Enc},\texttt{Dec})$ used for private channels must be IND-CPA secure.
    • NIZK Security: The NIZK system $(\textrm{Setup},\textrm{Prove},\textrm{Verify})$ must provide computational soundness and zero-knowledge.
    • Implicitly, security relies on the hardness of the Discrete Logarithm Problem (DLP) in the underlying elliptic curve group $\mathbb{G}$.
  • Security Definition: Security is defined using the Real/Ideal paradigm [31, 32]. We define an ideal functionality $\mathcal{F}_{FDKG}$ that captures the desired properties of FDKG (see Figure 3). The real protocol $\Pi_{FDKG}$ securely realizes $\mathcal{F}_{FDKG}$ if for any PPT adversary $\mathcal{A}$ controlling a set $\mathcal{C}$ in the real world, there exists a PPT simulator $\mathcal{S}$ controlling $\mathcal{C}$ in the ideal world such that no PPT environment $\mathcal{Z}$ can distinguish between interacting with $\mathcal{A}$ and the real protocol, or with $\mathcal{S}$ and the ideal functionality $\mathcal{F}_{FDKG}$.
V-B Ideal Functionality $\mathcal{F}_{FDKG}$
Ideal Functionality $\mathcal{F}^{n,t,k}_{FDKG}$
  • 1.
    Initialization:
    • Initializes an empty set of participants $\mathbb{D}=\emptyset$.
    • Initializes state flag ‘phase‘ = ’activation’.
  • 2.
    • If ‘phase‘ is ’activation’ and $P_{i}\notin\mathbb{D}$:
      • Add $P_{i}$ to $\mathbb{D}$.
      • Send (Activated, $i$) back to the simulator $\mathcal{S}$.
  • 3.
    • If ‘phase‘ is ’activation’ and $P_{i}\notin\mathbb{D}$:
      • Add $P_{i}$ to $\mathbb{D}$.
  • 4.
    Key Computation Trigger: Upon receiving (ComputeKeys) from the simulator $\mathcal{S}$ [^1] Note on Round Termination: For simplicity in the ideal functionality $\mathcal{F}_{FDKG}$, we model the transition from the activation phase to the key computation phase via this ‘(ComputeKeys)‘ trigger sent by the simulator. In a real-world deployment, particularly on a blockchain, the end of the distribution phase would typically be determined by external environmental factors like block height or a predetermined time duration, not controlled by the adversary. Participants submitting messages after this cutoff would be excluded. Our security analysis focuses on the correctness and privacy properties based on valid participation within the allowed timeframe, assuming the round termination mechanism itself is robust. The core security guarantees derived from the underlying cryptography (IND-CPA PKE, NIZK soundness/ZK) are not fundamentally affected by this modeling simplification.:
    • If ‘phase‘ is ’activation’:
      • Set ‘phase‘ = ’computed’.
V-C Security of the FDKG Protocol
We claim that our FDKG protocol ($\Pi_{FDKG}$, described in Section IV) securely realizes the ideal functionality $\mathcal{F}_{FDKG}$.
Theorem 1 (Security of FDKG Generation).
Assume the public-key encryption scheme $(\texttt{Enc},\texttt{Dec})$ used for the private channel is IND-CPA secure, and the NIZK $(\textrm{Setup},\textrm{Prove},\textrm{Verify})$ is zero-knowledge and computationally sound. Then, the protocol $\Pi_{FDKG}$ securely realizes the ideal functionality $\mathcal{F}_{FDKG}$ against static adversaries corrupting any set $\mathcal{C}\subset\mathbb{P}$.
Proof Sketch (Simulation Argument).
We demonstrate the security of $\Pi_{FDKG}$ by constructing a PPT simulator $\mathcal{S}$ that interacts with the ideal functionality $\mathcal{F}_{FDKG}$ (Figure 3) and any PPT adversary $\mathcal{A}$ attacking the real protocol $\Pi_{FDKG}$. The goal is to show that the view of $\mathcal{A}$ when interacting with $\mathcal{S}$ in the ideal world is computationally indistinguishable from its view when interacting with honest parties running $\Pi_{FDKG}$ in the real world. This implies that $\Pi_{FDKG}$ securely realizes $\mathcal{F}_{FDKG}$.
The simulator $\mathcal{S}$ internally simulates the execution environment for $\mathcal{A}$. It proceeds as follows:
@@ V-C Security of the FDKG Protocol
    Setup: $\mathcal{S}$ receives the corrupted set $\mathcal{C}$ from $\mathcal{A}$ and forwards it to $\mathcal{F}_{FDKG}$ via (Init, $\mathcal{C}$). $\mathcal{S}$ receives PKE public keys $\{\textrm{pk}_{i}\}_{i\in\mathcal{C}}$ from $\mathcal{A}$ and generates honest key pairs $(\textrm{pk}_{i},\textrm{sk}_{i})$ for $i\in\mathcal{H}$. $\mathcal{S}$ provides $\mathcal{A}$ with all public PKE keys $\{\textrm{pk}_{i}\}_{i\in\mathbb{P}}$.
@@
    Simulating Round 1 (Distribution Phase):
    • Corrupt Party Actions: When $\mathcal{A}$ decides a corrupt party $P_{i}\in\mathcal{C}$ should participate and broadcasts its message $(E_{i},\{C_{i,j}\}_{j\in\mathcal{G}_{i}},\pi_{FDKG_{i}})$:
      • $\mathcal{S}$ verifies the proof $\pi_{FDKG_{i}}$ using the public parameters and statement components derived from the message. If invalid, $\mathcal{S}$ flags this internally.
      • $\mathcal{S}$ extracts the claimed polynomial $f_{i}$ and guardian set $\mathcal{G}_{i}$ from $\mathcal{A}$’s internal state or message.
      • $\mathcal{S}$ sends (ActivateCorrupt, $i,f_{i},\mathcal{G}_{i}$) to $\mathcal{F}_{FDKG}$.
    • Simulating Honest Party Actions: $\mathcal{S}$ must decide which honest parties $P_{i}\in\mathcal{H}$ participate in this round. This decision can be based on instructions from the environment or a default strategy (e.g., all honest parties participate). For each honest party $P_{i}$ chosen to participate:
      • $\mathcal{S}$ selects a guardian set $\mathcal{G}_{i}$ for $P_{i}$ (consistent with honest behavior, e.g., randomly chosen from $\mathbb{P}\setminus\{P_{i}\}$).
      • $\mathcal{S}$ informs the ideal functionality by sending (Activate, $\mathcal{G}_{i}$) on behalf of $P_{i}$. $\mathcal{S}$ subsequently receives confirmation (Activated, $i$) from $\mathcal{F}_{FDKG}$.
    • Triggering Key Computation in Ideal World: Once $\mathcal{S}$ determines that all parties wishing to participate (from $\mathcal{A}$’s perspective) have declared their intent (i.e., corrupt parties have sent messages, honest parties’ participation decided by $\mathcal{S}$), $\mathcal{S}$ sends (ComputeKeys) to $\mathcal{F}_{FDKG}$.
    • Receiving Ideal Outputs: $\mathcal{S}$ receives the message (KeysComputed, E, $\{E_{k}\}_{k\in\mathbb{D}}$, $\{\mathcal{G}_{k}\}_{k\in\mathbb{D}\cap\mathcal{H}}$, corrupt party details, $\{[d_{k}]_{j}\mid k\in\mathbb{D},P_{j}\in\mathcal{G}_{k}\cap\mathcal{C}\}$) from $\mathcal{F}_{FDKG}$. This defines the set $\mathbb{D}$ of valid participants whose contributions were accepted by the ideal functionality.
    • Constructing and Delivering Simulated Broadcasts: For each honest party $P_{i}\in\mathbb{D}\cap\mathcal{H}$, $\mathcal{S}$ constructs a simulated broadcast message $(P_{i},E_{i},\mathcal{G}_{i},\{C_{i,j}\}_{P_{j}\in\mathcal{G}_{i}},\pi_{FDKG_{i}}^{sim})$ where:
      • $E_{i}$ and $\mathcal{G}_{i}$ are taken from the information received from $\mathcal{F}_{FDKG}$.
      • $\{C_{i,j}\}_{P_{j}\in\mathcal{G}_{i}}$ consists of ciphertexts encrypting dummy values (e.g., 0). For each $P_{j}\in\mathcal{G}_{i}$, $\mathcal{S}$ samples PKE randomness $\textrm{rand}_{i,j}$ and computes $C_{i,j}\leftarrow\texttt{Enc}(\textrm{pk}_{j},0,\textrm{rand}_{i,j})$.
      • $\pi_{FDKG_{i}}^{sim}$ is a simulated NIZK proof generated using the NIZK simulator: $\pi_{FDKG_{i}}^{sim}\leftarrow\textrm{Sim}_{\textrm{NIZK}}(\textrm{crs},(E_{i},\{\textrm{pk}_{j}\}_{P_{j}\in\mathcal{G}_{i}},\{C_{i,j}\}_{P_{j}\in\mathcal{G}_{i}}))$.
      $\mathcal{S}$ then presents the complete set of broadcast messages (real ones from $\mathcal{A}$ for corrupt $i\in\mathbb{D}\cap\mathcal{C}$, and simulated ones $M_{i}^{sim}$ for honest $i\in\mathbb{D}\cap\mathcal{H}$) to $\mathcal{A}$.
@@
    Final Output: $\mathcal{S}$ runs $\mathcal{A}$ until it terminates and outputs whatever $\mathcal{A}$ outputs.
Indistinguishability Argument: The core of the security proof is arguing that the view provided to $\mathcal{A}$ by $\mathcal{S}$ in the ideal world is computationally indistinguishable from $\mathcal{A}$’s view in the real world. This relies on the security properties of the underlying cryptographic primitives:
  • Simulated vs. Real NIZK Proofs: The proofs $\pi_{FDKG_{i}}$ generated by honest parties in the real world are replaced by simulated proofs $\pi_{FDKG_{i}}^{sim}$ generated by $\mathcal{S}$ using the NIZK simulator. The zero-knowledge property of the NIZK system guarantees that $\mathcal{A}$ cannot distinguish between real and simulated proofs.
  • Simulated vs. Real Encrypted Shares: Consider the ciphertexts $\{C_{i,j}\}$ broadcast by an honest party $P_{i}$.
    • If the recipient $P_{j}$ is also honest ($P_{j}\in\mathcal{H}$), $\mathcal{A}$ does not possess the secret key $\textrm{sk}_{j}$ required for decryption. In the real world, $C_{i,j}$ encrypts the real share $s_{i,j}$. In the simulation, it encrypts a dummy value (e.g., 0). By the IND-CPA security of the PKE scheme, $\mathcal{A}$ cannot distinguish these ciphertexts.
    • If the recipient $P_{j}$ is corrupt ($P_{j}\in\mathcal{C}$), $\mathcal{A}$ possesses $\textrm{sk}_{j}$. In the real world, $C_{i,j}$ encrypts the real share $s_{i,j}$. In the simulation, $\mathcal{F}_{FDKG}$ provides the actual share $s_{i,j}$ to $\mathcal{S}$. $\mathcal{S}$ can use this $s_{i,j}$ to compute the ”correct” ciphertext $C_{i,j}$ using fresh randomness and provide this to $\mathcal{A}$ in the simulated broadcast message $M_{i}^{sim}$. Alternatively, if $\mathcal{S}$ encrypts a dummy value, the indistinguishability relies on the combination of IND-CPA and the fact that the overall view remains consistent due to the later steps. The crucial aspect is that $\mathcal{S}$ can provide $\mathcal{A}$ with exactly the information it is entitled to (shares sent to corrupted parties).
  • Consistency of Public Keys: The partial public keys $E_{i}$ for honest participants are provided by $\mathcal{F}_{FDKG}$ based on its internal random polynomial choices. For corrupt participants, $\mathcal{A}$ provides $E_{i}$ along with $\pi_{FDKG_{i}}$. The computational soundness of the NIZK ensures that $\mathcal{A}$ cannot produce a valid proof $\pi_{FDKG_{i}}$ for a pair $(E_{i},\mathbb{C}_{i})$ that is inconsistent with some underlying polynomial $f_{i}(X)$ (up to negligible probability). Therefore, the set of partial keys $\{E_{i}\}_{i\in\mathbb{D}}$ and the derived global key E have computationally indistinguishable distributions in the real and ideal worlds.
  • Participation and Guardian Sets: The process by which the set of valid participants $\mathbb{D}$ is determined (based on valid proofs) and how guardian sets $\mathcal{G}_{i}$ for honest parties are chosen and communicated is consistent between the simulation and the real protocol execution flow.
Combining these points, we conclude that any PPT adversary $\mathcal{A}$’s view in the simulation is computationally indistinguishable from its view in the real protocol execution, assuming the underlying PKE scheme is IND-CPA secure and the NIZK system is zero-knowledge and computationally sound. This completes the proof sketch. ∎
V-D FDKG Security Properties
The simulation-based security theorem (Theorem 1) implies the following essential security properties for the FDKG generation phase (Round 1):
  • Correctness: For the set of valid participants $\mathbb{D}$, the protocol outputs are correctly formed according to the protocol specification. That is, the computed global public key E corresponds to the sum of the implicit partial secret keys $\textbf{E}=G^{\sum_{i\in\mathbb{D}}d_{i}}$, each partial public key relates correctly to its partial secret key $E_{i}=G^{d_{i}}$, and the encrypted shares $C_{i,j}$ correspond to the correct evaluations $f_{i}(j)$ of the underlying polynomial $f_{i}(X)$ defining $d_{i}=f_{i}(0)$. This is guaranteed by the definition of $\mathcal{F}_{FDKG}$ in the ideal world and by the computational soundness of the NIZK proofs $\pi_{FDKG_{i}}$ in the real world.
  • Privacy: The adversary $\mathcal{A}$ (controlling the set $\mathcal{C}$) learns no information about the partial secret keys $d_{i}$ of honest participants ($i\in\mathcal{H}\cap\mathbb{D}$) or the final global secret key d, beyond what can be inferred from the public information (E, $\{E_{i}\}_{i\in\mathbb{D}}$) and the shares explicitly sent to corrupted parties ($\{[d_{i}]_{j}\mid P_{j}\in\mathcal{G}_{i}\cap\mathcal{C}\}$). This follows from the indistinguishability argument in the proof sketch: the simulator $\mathcal{S}$ constructs $\mathcal{A}$’s view without access to honest parties’ secrets, relying on the IND-CPA security of the PKE and the zero-knowledge property of the NIZK system.
  • Robustness (During Generation): Honest participating parties ($\mathcal{H}\cap\mathbb{D}$) successfully complete the key generation phase and obtain their outputs (their partial secret $d_{i}$, the global public key E, etc.) as defined by the ideal functionality $\mathcal{F}_{FDKG}$, irrespective of the (mis)behavior of corrupted parties. Malicious parties $P_{j}\in\mathcal{C}$ cannot prevent protocol completion for honest parties other than by failing to participate or by broadcasting messages with invalid proofs $\pi_{FDKG_{j}}$, which are simply ignored by observers (as reflected in the Post-Round 1 processing).
V-E Liveness of Key Reconstruction
Distinct from the security properties of the generation phase, Liveness addresses the success of the reconstruction phase (Round 2). It guarantees that the global secret key d can be recovered if enough participants and their guardians are available and behave honestly during this second phase. This property assumes the generation phase successfully realized $\mathcal{F}_{FDKG}$.
Definition 1 (Liveness of Reconstruction).
The FDKG protocol achieves Liveness of Reconstruction if, for every participant $P_{i}\in\mathbb{D}$ whose contribution $d_{i}$ is required for an application using the key, at least one of the following conditions holds during the reconstruction phase:
  • $P_{i}$ is honest and available, and successfully broadcasts $(P_{i},\text{secret},d_{i},\pi_{PS_{i}})$ where $\pi_{PS_{i}}$ is valid.
  • At least $t$ distinct parties $\{P_{j_{1}},\dots,P_{j_{t}}\}\subseteq\mathcal{G}_{i}$ are honest and available, and each successfully broadcasts $(P_{j_{r}},\text{share},P_{i},s_{i,j_{r}},\pi_{SPS_{i,j_{r}}})$ where $\pi_{SPS_{i,j_{r}}}$ is valid.
Liveness fails for a specific $d_{i}$ if the adversary controls $P_{i}$ (preventing the first condition) and controls at least kt+1k-t+1 parties in $\mathcal{G}_{i}$ (preventing the second condition) during the reconstruction phase.
Theorem 2 (Liveness of Reconstruction).
Assume the FDKG generation phase completed successfully (realizing $\mathcal{F}_{FDKG}$). Assume the NIZK proofs used in the reconstruction phase ($\pi_{PS_{i}}$, $\pi_{SPS_{j,i}}$) are computationally sound. The FDKG protocol achieves Liveness of Reconstruction if, for every participant $P_{i}\in\mathbb{D}$, the adversary does not control both $P_{i}$ and kt+1k-t+1 guardians in $\mathcal{G}_{i}$ during the reconstruction phase.
@@ V-E Liveness of Key Reconstruction
The proof relies on the properties of SSS and the soundness of the NIZK proofs $\pi_{PS_{i}}$ and $\pi_{SPS_{j,i}}$. For any $P_{i}\in\mathbb{D}$:
  1. 1.
    If $P_{i}$ is honest/available in Round 2, it broadcasts $d_{i}$ and $\pi_{PS_{i}}$. Soundness ensures $\pi_{PS_{i}}$ only verifies if $d_{i}$ corresponds to $E_{i}$.
  2. 2.
    If $P_{i}$ is corrupt/unavailable, reconstruction depends on its guardians $\mathcal{G}_{i}$. If the adversary controls fewer than kt+1k-t+1 guardians in $\mathcal{G}_{i}$, then at least $t$ guardians are honest/available. Each such honest guardian $P_{j}\in\mathcal{G}_{i}$ will decrypt $C_{i,j}$ to get $s_{i,j}$ using $\textrm{sk}_{j}$, and broadcast $s_{i,j}$ with $\pi_{SPS_{i,j}}$. Soundness ensures $\pi_{SPS_{i,j}}$ only verifies if $s_{i,j}$ is the correct decryption of $C_{i,j}$ using $\textrm{sk}_{j}$. Correctness of Round 1 (guaranteed by soundness of $\pi_{FDKG_{i}}$) ensures $C_{i,j}$ encrypted $f_{i}(j)$. Thus, the observer obtains at least $t$ valid shares $\{s_{i,j}=f_{i}(j)\}$. By SSS properties, these $t$ shares allow reconstruction of $d_{i}=f_{i}(0)$.
The theorem’s condition ensures that at least one of these two recovery methods is available for each $d_{i}$, thus guaranteeing $\textbf{d}=\sum d_{i}$ can be computed. ∎

### v3

@@ V Security Analysis
This section analyzes the security of $\Pi_{FDKG}$ in two phases: Generation (Round 1) and Reconstruction (Round 2). We adopt the standard real/ideal paradigm [49, 50]. A detailed ideal functionality $\mathcal{F}_{FDKG}$ appears in Appendix A.
V-A Security Model and Assumptions
@@
    Parties and communication. There are $n$ potential parties $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$. Communication occurs over an authenticated public broadcast channel in synchronous rounds.
@@
    Adversary. Unless stated otherwise, the adversary is static probabilistic polynomial time (PPT) and corrupts a fixed subset $\mathcal{C}\subset\mathbb{P}$ before the protocol starts; corrupted parties are fully Byzantine. Let $\mathcal{H}=\mathbb{P}\setminus\mathcal{C}$ denote the honest set. Honest parties may be unavailable in Round 2 (crash/absence), which affects liveness. Section VI-C discusses adaptive adversaries and targeted attacks.
@@
    Cryptographic assumptions.
    • PKE security. The channel for encrypted shares $(\textrm{Enc},\textrm{Dec})$ (Section III-F) is assumed IND-CPA secure. In our implementation this is an ElGamal variant over the BabyJub subgroup. Security reduces to the DDH assumption (and ultimately DLP hardness) in this group; ciphertext malleability is acceptable because correctness is enforced by NIZK soundness. Any alternative IND-CPA PKE could be substituted.
    • NIZK security. The proof system $(\textrm{Setup},\textrm{Prove},\textrm{Verify})$ (Section III-E) is assumed computationally sound and zero-knowledge in the CRS model. We instantiate Groth16 over the BN254 pairing curve (also known as bn128, the default in circom/snarkjs); this uses a circuit-specific structured reference string (SRS) generated in a trusted setup. Soundness relies on standard pairing-group assumptions on BN254; zero-knowledge follows from the Groth16 simulator. Other systems (e.g., PLONK/Marlin with a universal updatable SRS [43, 44], or STARKs with transparent setup [45]) are compatible at the level of assumptions used here, trading proof size/proving time for setup properties and security level.
    • Discrete logarithm hardness. We assume DLP hardness in (i) the BabyJub subgroup used for ElGamal and (ii) the pairing groups induced by the chosen SNARK curve, as applicable.
V-B Security of the Generation Phase (Round 1)
We prove Generation-phase security in the real/ideal paradigm using the ideal functionality $\mathcal{F}_{FDKG}$ (Appendix A). Informally, the simulator replaces honest encrypted shares by encryptions of 0 and uses the NIZK simulator to generate proofs; IND-CPA and zero-knowledge imply indistinguishability, while NIZK soundness enforces well-formedness. The following theorem summarizes the resulting properties.
Theorem 1 (Generation: Correctness, Privacy, Robustness).
Assume the PKE is IND-CPA and the NIZK is zero-knowledge and computationally sound. For the set of valid participants $\mathbb{D}$ determined by verification:
  1. 1.
    Correctness. For each $i\in\mathbb{D}$, $E_{i}=G^{d_{i}}$ and
    $\textbf{E}\;=\;\prod_{i\in\mathbb{D}}E_{i}\;=\;G^{\sum_{i\in\mathbb{D}}d_{i}}.$
    Moreover, every published $C_{i,j}$ encrypts $f_{i}(j)$ for a degree-(t1)(t\!-\!1) polynomial $f_{i}$ with $f_{i}(0)=d_{i}$.
  2. 2.
    Privacy (during Generation). The adversary’s view leaks no information about $\{d_{i}\}_{i\in\mathbb{D}\cap(\mathbb{P}\setminus\mathcal{C})}$ beyond $(\textbf{E},\{E_{i}\}_{i\in\mathbb{D}})$ and the shares explicitly addressed to corrupted recipients $\{[d_{i}]_{j}\mid j\in\mathcal{G}_{i}\cap\mathcal{C}\}$.
  3. 3.
    Robustness (completion). Malformed broadcasts (invalid proofs) are rejected; corrupted parties cannot force incorrect acceptance. Honest parties with valid broadcasts are included in $\mathbb{D}$ and obtain their outputs as per $\mathcal{F}_{FDKG}$.
@@ V-B Security of the Generation Phase (Round 1)
Simulate honest parties by encrypting 0 and using simulated NIZKs; IND-CPA and zero-knowledge yield indistinguishability. NIZK soundness binds each dealer’s ciphertexts to a single polynomial and $E_{i}$, ensuring well-formedness. Verification excludes malformed data, implying robustness. ∎
V-C Security of the Reconstruction Phase (Round 2)
This subsection assumes the Generation phase completed successfully (Theorem 1) and that the Round 2 NIZKs ($\pi_{PS_{i}}$, $\pi_{SPS_{i,j}}$) are computationally sound. We analyze Correctness, Privacy (reconstruction), and Liveness.
Notation for Reconstruction
Define the family of reconstruction-capable sets
$R\;:=\;\Bigl\{\,S\subseteq\mathbb{P}\ \Big|\ \forall i\in\mathbb{D}:\ i\in S\ \ \lor\ \ |S\cap\mathcal{G}_{i}|\geq t\,\Bigr\}.$ (1)
Intuitively, $S\in R$ iff, for every participant $i$, $S$ contains either $i$ or at least $t$ of $i$’s guardians. Round 2 privacy and liveness depend on the guardian topology $\{\mathcal{G}_{i}\}$ via $R$ rather than on a single global corruption fraction, as in traditional uniform DKGs.
Theorem 2 (Reconstruction: Correctness).
If for every $i\in\mathbb{D}$ either (i) there is a valid broadcast $(P_{i},\text{secret},d_{i},\pi_{PS_{i}})$, or (ii) there exist $t$ distinct guardians $j\in\mathcal{G}_{i}$ with valid broadcasts $(P_{j},\text{share},P_{i},s_{i,j},\pi_{SPS_{i,j}})$, then any observer reconstructs the unique $\textbf{d}=\sum_{i\in\mathbb{D}}d_{i}.$ Moreover, checking $G^{d_{i}}=E_{i}$ prevents substitution.
Proof sketch.
By SSS, any $t$ correct shares interpolate $f_{i}(0)=d_{i}$. Soundness of $\pi_{PS_{i}}$ and $\pi_{SPS_{i,j}}$ ensures that only values consistent with Round-1 commitments are accepted. Summing over $i$ yields d. ∎
Theorem 3 (Reconstruction: Privacy).
The adversary can reconstruct the global secret if and only if there exists a set $S\subseteq\mathcal{C}$ such that $S\in R$. Equivalently, reconstruction privacy holds iff $\mathcal{C}\notin R$.
Proof sketch.
If the adversary reconstructs, it must possess, for each $i$, either $i$’s partial secret (when $i\in\mathcal{C}$) or $t$ decryptable guardian shares (when $t$ guardians are corrupted). Collect those principals into $S$; then $S\subseteq\mathcal{C}$ and $S\in R$.
Conversely, if $S\subseteq\mathcal{C}$ and $S\in R$, then for every $i$ the adversary controls either $i$ or $t$ guardians. In the first case it knows $d_{i}$; in the second it decrypts $t$ shares and reconstructs $d_{i}$. Summing yields d. ∎
Theorem 4 (Reconstruction: Liveness).
The adversary cannot prevent reconstruction if and only if there exists a set $S\in R$ such that $S\subseteq U\setminus\mathcal{C}$. Equivalently,
$\forall i\in\mathbb{D}:\ \bigl(i\notin\mathcal{C}\bigr)\ \ \lor\ \ \bigl(|\mathcal{C}\cap\mathcal{G}_{i}|\leq k-t\bigr).$
Proof sketch.
If such $S$ exists, then $S$ contains, for each $i$, either $i$ or $t$ honest guardians; reconstruction proceeds as in Theorem 2.
Conversely, if no such $S$ exists, then for some $i$ the adversary controls $i$ and at least kt+1k-t+1 of $\mathcal{G}_{i}$, preventing both direct revelation and a $t$-guardian reconstruction path; hence liveness fails. ∎
Potential attacks and mitigations
Share withholding / selective opening in Round 2 is captured by Theorem 4: liveness fails for $i$ only if the adversary both corrupts $i$ and at least kt+1k{-}t{+}1 of $\mathcal{G}_{i}$. Forgery or equivocation is ruled out by soundness of $\pi_{PS_{i}}$ and $\pi_{SPS_{i,j}}$. Guardian concentration attacks (attracting many delegations to a small set of guardians) increase redundancy under random churn but reduce privacy thresholds under adaptivity (Theorem 3); deployers can mitigate by (i) capping per-guardian incoming delegations, (ii) diversifying recommendation lists, (iii) sampling a fraction of guardians via a verifiable random function (VRF), and (iv) rate-limiting or staking for high-degree guardians. Finally, Sybil guardians are out of scope of our PKI model but can be addressed operationally via admission controls (identity or stake).
On topology dependence
Two extreme cases illustrate the role of overlap:
  • Disjoint guardians. If $\mathcal{G}_{i}\cap\mathcal{G}_{i^{\prime}}=\varnothing$ for $i\neq i^{\prime}$, then making a fixed $i$ unrecoverable requires at least 1+(kt+1)1+(k-t+1) targeted corruptions (participant $i$ and kt+1k-t+1 distinct guardians).
  • Full overlap. If $\mathcal{G}_{i}=\mathcal{G}$ for all $i$, then corrupting kt+1k-t+1 parties in $\mathcal{G}$ makes every corrupted participant $i$ unrecoverable. Thus, the fraction of corruptions needed to break liveness or privacy depends on $\{\mathcal{G}_{i}\}$.
Next section empirically studies overlap policies (Random vs. Barabási–Albert) and their impact on success rates.